a science 3. Simmons leadership team does some research and learns that if it doesnt build building, not only will it not bAnswer question #4.
Answer #1

Q1 Answer

Existing funds $          6,55,00,000 Given
Debt Needed $           3,45,00,000 Given
Total to invest in new building $        10,00,00,000 Given
Interest rate on debt 4.50% Given
cost of debt $             15,52,500 Interest rate on debt * Debt Needed
New joinees per year 100 Given
Tution per year $                      37,500 Given
Annual income per year $             37,50,000 New joinees per year * Tution per year
Annual profit /loss $             21,97,500 Annual income per year - cost of debt

Cost of debt would be annual interest we need to pay on the amount of debt taken that is Rate of interest * amount of debt

Annual income per year is how much we would earn from the amount we would invest that would be from the Tution fees from the 100 new students that would be joining us = Number of new Joinee * Tution fees per student

Annual profilt /oss would be difference of the amount we have invested and the income we are getting from that amount =

annual income - Cost of Debt

Q 2 =

Amount to invest $   6,55,00,000.00 Given
investment annual growth rate 6.75% Given
Annual income per year $     44,21,250 Amount to invest * investment annual growth rate
Tution per year $              37,500.00 Given
Number of whole scholarship that could be funded 118 Annual interest income per year / tution per year

Here we are planning to invest the amount and not take any debt. If we choose to invest that amount at some interest rate(given) then we would earn some interest from that amount which could be used to fund scholarship for certain students.

Hence - Annal income = this would be the Amount to invest * investment annual growth rate

Number of scholarship could be funded from the income we get = Annual interest income per year / tution per year ( rounding off to nearest decimal)

Q3 a )

In case simmons doesnt invest in the new building then it would loose some students = 100

Tution per year $                      37,500
Number of graduate student 100
Revenue from 100 students $             37,50,000

Hence the revenue lost would be - tution fees collected from 1 student * number of graduate students lost = - $3750000

(negative because its a loss)

Q3 b) %age of that money that this loss represents would be Amount to invest / revenue lost = 65500000 / 3750000 = 5.73 %

Q 4) Considering all 3 scenarios at once

that is - if simmons doesnt invest it would have a loss of - $3750000 (from part c),

BUT it could invest the amount of 65500000 at 6.75 % and earn 4421250 which could be used for funding 118 new students (from part b)

Hence It would loose 100 students BUT could give scholarship to 118 students = +118 - 100 = 18

Still get 18 more students every year

Profit = 18 * 37500 = 675000

SO IT SHOULD INVEST THE MONEY AT 6.75 % AND USE THAT MONEY TO GIVE SCHOLARSHIP TO NEW STUDENTS IF POSSIBLE

ELSE IT SHOULD GO WITH OPTION A = THAT IS BUILD A NEW BUILDING BY CREATING DEBT AND GET PROFIT OF       21,97,500 PER YEAR.